# FIARSE: Model-Heterogeneous Federated Learning via Importance-Aware Submodel Extraction

### v1

@@ C.2 Preliminary
$\leq\frac{2A}{N}\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\mathbb{E}\left\|\nabla F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (44)
@@
$\quad+\frac{2A}{NK}\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (45)
@@
$=\frac{2A}{N}\sum_{i\in[N]}\mathbb{E}\left\|\nabla F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (46)
@@
$\quad+\frac{2A}{N}\sum_{i\in[N]}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[0:\gamma_{i}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[0:\gamma_{i}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (47)
@@
$\leq 2A\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+72KA\eta_{l}^{2}L^{2}\left(\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+\sigma^{2}+\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)$ (49)
where Equation (46) and Equation (47) are applied with Assumption 4.3 and Assumption 4.1, respectively, to attain Equation (48). Finally, we apply with Lemma C.2 and attain Equation (49). With the learning rate set as described, we can obtain the desired conclusion. ∎
@@
Suppose that Assumption 4.1, 4.2 and 4.3 hold. Let the local learning rate satisfy $\eta_{l}\leq\min\left(\frac{1}{2L\sqrt{K(K+1)}},\frac{1}{6L\sqrt{(K+1)A}}\right)$. With FIARSE, we have the following conclusion:
@@
Knowing that the recursive function of the global updates of FIARSE follows that
@@
$\tilde{\boldsymbol{x}}_{t+1}-\tilde{\boldsymbol{x}}_{t}=-\eta_{s}\eta_{l}\textsf{Agg}_{i\in\mathcal{A}}\left(\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right),$ (51)
we have
$\quad\mathbb{E}\left\|\tilde{\boldsymbol{x}}_{t+1}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}=\mathbb{E}\sum_{j=0}^{n}\left\|\frac{\eta_{s}\eta_{l}}{\left|\mathcal{A}_{\gamma_{j}^{\prime}}\right|}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (52)
$\leq\eta_{s}^{2}\eta_{l}^{2}\cdot\mathbb{E}\sum_{j=0}^{n}\frac{K}{\left|\mathcal{A}_{\gamma_{j}^{\prime}}\right|}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (53)
$=\eta_{s}^{2}\eta_{l}^{2}K\cdot\sum_{j=0}^{n}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (54)
Next, let us analyze the bound for $\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$:
@@
$\quad\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (55)
@@
$\leq 4\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)\right\|_{2}^{2}$ (56)
@@
$\quad+4\mathbb{E}\left\|\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right)-\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}$ (57)
$\quad+4\mathbb{E}\left\|\nabla_{\tilde{\boldsymbol{x}}_{t}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\odot\mathcal{M}^{(i)}_{t}\left(\tilde{\boldsymbol{x}}_{t}\right)-\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}+4\mathbb{E}\left\|\nabla F^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}$ (58)
Following the analysis mentioned in Lemma C.3, we have:
@@
$\quad\sum_{j=0}^{n}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (59)
$\leq\frac{4A}{N}\sum_{i\in[N]}\sum_{k=0}^{K-1}L^{2}\mathbb{E}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+\frac{4KA}{N}\sum_{i\in[N]}\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+4K\sigma^{2}+4K\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}$ (60)
By Lemma C.2, we can further simplify the inequality for
@@
$\quad\sum_{j=0}^{n}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right\|_{2}^{2}$ (61)
$\leq 4KA\left(1+36\eta_{l}^{2}L^{2}(1+K)\right)\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+4K\sigma^{2}\left(1+36\eta_{l}^{2}L^{2}(1+K)\right)$

### v2

@@ C.2 Preliminary
$\mathbb{E}\left\|\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\right\|_{2}^{2}$ $\leq 2\mathbb{E}\left\|\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}\hat{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}+2\mathbb{E}\left\|\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}$ (52)
@@
$\leq 2\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}\mathbb{E}\left\|\hat{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}+2\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}$ (53)
With the inequality above, the LHS of Equation (48) is bounded by
@@
$\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\right\|_{2}^{2}\leq$ $2\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}\left|N_{\gamma_{j}^{\prime}}\right|}\sum_{i\in N_{\gamma_{j}^{\prime}}}\mathbb{E}\left\|\hat{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}$ (54)
@@
$+2\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}\left|N_{\gamma_{j}^{\prime}}\right|K}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}$ (55)
According to Lemma C.2, we have
$\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}\left|N_{\gamma_{j}^{\prime}}\right|}\leq\frac{A}{N}.$ (56)
Therefore, we further bound Equation (55) for
$\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\right\|_{2}^{2}$ (57)
$\leq$ $\frac{2A}{N}\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\mathbb{E}\left\|\hat{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t}\right\|_{2}^{2}+\frac{2A}{NK}\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}_{t,k}\right\|_{2}^{2}$ (58)
$=$ $\frac{2A}{N}\sum_{i\in[N]}\mathbb{E}\left\|\hat{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t}\right\|_{2}^{2}+\frac{2A}{NK}\sum_{i\in[N]}\sum_{k=0}^{K-1}\mathbb{E}\left\|\tilde{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t,k}\right\|_{2}^{2}$ (59)
where the last equation holds because the second norm is non-zero when the client $i\in[N]$ owns the part $[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}],j\in\{0,\dots,n\}$, and different parts (i.e., $[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]$ for different $j$s) are independent with each other when calculating the second norm.
With Assumption 5.3 and Assumption 5.1, we separately bound two terms in Equation (59) for
$\left\|\hat{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t}\right\|_{2}^{2}\leq\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2};\qquad\left\|\tilde{\boldsymbol{g}}^{[0:\gamma_{i}](i)}_{t,k}\right\|_{2}^{2}\leq L^{2}\left\|\boldsymbol{x}_{t,k}^{(i)}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}.$ (60)
Therefore, With the inequality above, the LHS of Equation (48) is bounded by
$\sum_{j=0}^{n}\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{v}_{t}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\right\|_{2}^{2}$ (61)
@@
$\leq$ $2A\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+72KA\eta_{l}^{2}L^{2}\left(\delta_{t}^{2}\left\|\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}+\sigma^{2}+\left\|\nabla F\left(\tilde{\boldsymbol{x}}_{t}\right)\right\|_{2}^{2}\right)$ (63)
Since the learning rate meets the constraint of Lemma C.3, Equation (63) is obtained when Lemma C.3 applies. Furthermore, with the defined learning rate, we can attain the desired conclusion of Equation (48). ∎
@@
Suppose that Assumption 5.1, 5.2 and 5.3 hold. Let the local learning rate satisfy $\eta_{l}\leq\min\left(\frac{1}{2L\sqrt{K(K+1)}},\frac{1}{6L\sqrt{(K+1)A}}\right)$. With FIARSE, we have the following conclusion:
@@
The recursive function of the global updates of FIARSE follows that
@@
$\tilde{\boldsymbol{x}}_{t+1}-\tilde{\boldsymbol{x}}_{t}=-\eta_{s}\eta_{l}\textsf{Agg}_{i\in\mathcal{A}}\left(\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right).$ (65)
Let us define $\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}$ to be
@@
$\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}=\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}]}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right).$ (66)
Therefore,
@@
$\textsf{Agg}_{i\in\mathcal{A}}\left(\sum_{k=0}^{K-1}\nabla_{\boldsymbol{x}_{t,k}^{(i)}}F_{i}\left(\boldsymbol{x}_{t,k}^{(i)}\odot\mathcal{M}^{(i)}_{t}\left(\boldsymbol{x}_{t,k}^{(i)}\right)\right)\right)=\bigcup_{j\in[n]}\frac{1}{\left|\mathcal{A}_{\gamma_{j}^{\prime}}\right|}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}.$ (67)
The bound for $\mathbb{E}\left\|\tilde{\boldsymbol{x}}_{t+1}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}$ is formulated and simplified as follows:
@@
$\mathbb{E}\left\|\tilde{\boldsymbol{x}}_{t+1}-\tilde{\boldsymbol{x}}_{t}\right\|_{2}^{2}=$ $\mathbb{E}\sum_{j=0}^{n}\left\|\frac{\eta_{s}\eta_{l}}{\left|\mathcal{A}_{\gamma_{j}^{\prime}}\right|}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (68)
@@
$\leq$ $\eta_{s}^{2}\eta_{l}^{2}\cdot\mathbb{E}\sum_{j=0}^{n}\frac{K}{\left|\mathcal{A}_{\gamma_{j}^{\prime}}\right|}\sum_{i\in\mathcal{A}_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\left\|\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (69)
@@
$=$ $\eta_{s}^{2}\eta_{l}^{2}K\cdot\sum_{j=0}^{n}\sum_{i\in N_{\gamma_{j}^{\prime}}}\sum_{k=0}^{K-1}\frac{1}{\left|N_{\gamma_{j}^{\prime}}\right|}\cdot\frac{\binom{N}{A}-\binom{N-\left|N_{\gamma_{j}^{\prime}}\right|}{A}}{\binom{N}{A}}\cdot\mathbb{E}\left\|\boldsymbol{g}_{t,k}^{[\gamma_{j-1}^{\prime}:\gamma_{j}^{\prime}](i)}\right\|_{2}^{2}$ (70)
Equation (69) is obtained based on Cauchy–Shwarz inequality, and Equation (70) holds according to Lemma C.1.
Let us define